OpenStudy (anonymous):

Someone pushes a 15 kg object with a force of 500N for a distance of 1 meter. The object started from rest and 2 meters from the block is a spring with k = 20 N/m. The object slides on a surface with coefficient of friction .2. How far is the spring compressed by the object Someone pushes a 15 kg object with a force of 500N for a distance of 1 meter. The object started from rest and 2 meters from the block is a spring with k = 20 N/m. The object slides on a surface with coefficient of friction .2. How far is the spring compressed by the object @Mathematics

OpenStudy (anonymous):

Fnet = Fapp - Ff Fnet = 500 - (150)(.2) Fnet =470 N F = ma 470 = 15 a a= 470/15 a= 31.33333 Vf = (2(31.33333)1)^.5 Vf = 7.91622764 .5mv^2 = .5kx^2 mv^2 = kx^2 15(7.91622764)^2 = 20x^2 x = ((15*(7.91622764) ^2) / 20 )^.5 x = 6.85565424m Hope it works.. =)